# Analysis of a multi-target linear shrinkage covariance estimator

### v1

@@ 3 Oracle and proposed estimator
$\bar{b}_{T^{(i)},n}^{2}=\frac{p_{n}}{(n-1)^{2}}\sum_{k=1}^{n}\left\langle\tilde{X}_{\cdot k}^{n}\left(\tilde{X}_{\cdot k}^{n}\right)^{T},T_{n}^{(i)}\right\rangle_{n}^{2}-\frac{1}{n}\left\langle S_{n}T_{n}^{(i)},T_{n}^{(i)}S_{n}\right\rangle_{n},$ (12)
$\hat{\mathbb{V}}\left(\langle S_{n},T_{n}^{(i)}\rangle_{n}\right)=q_{n}^{(0)}\bar{b}_{T^{(i)},n}^{2}+q_{n}^{(1)}\left\langle S_{n}T_{n}^{(i)},T_{n}^{(i)}S_{n}\right\rangle_{n}-q_{n}^{(2)}\langle S_{n},T_{n}^{(i)}\rangle_{n}^{2},$
$q_{n}^{(0)}=\frac{(n-1)^{2}}{p_{n}(n-2)(n-3)},q_{n}^{(1)}=\frac{n-1}{p_{n}n(n-2)},q_{n}^{(2)}=\frac{n^{2}-2n-1}{n(n-2)(n-3)}.$
@@
The proof works around the fact that $\left(T_{n}^{(k)}\right)_{k}$ is an orthonormal family to use Assumption 2 then.
Theorem 1 (Variance estimators).
Denoting $\mathbb{V}(S_{n})=\mathbb{E}[\lVert S_{n}-\Sigma_{n}\rVert_{n}^{2}]$ and $\mathbb{V}\left(\langle S_{n},T_{n}^{(i)}\rangle_{n}\right)=\mathbb{E}\left[\langle S_{n}-\Sigma_{n},T_{n}^{(i)}\rangle_{n}^{2}\right]$, under assumptions 1 to 4, in both situations where the mean is known or unknown for all $i\in\llbracket 1,N_{n}\rrbracket$, $\hat{\mathbb{V}}(S_{n})$ and $\hat{\mathbb{V}}(\langle S_{n},T_{n}^{(i)}\rangle_{n})$ are unbiased estimators of $\mathbb{V}(S_{n})$ and $\mathbb{V}(\langle S_{n},T_{n}^{(i)}\rangle_{n})$ respectively. Moreover, $\mathbb{E}\left[\left(\hat{\mathbb{V}}(S_{n})-\mathbb{V}(S_{n})\right)^{2}\right]\longrightarrow 0$ and $\mathbb{E}\left[\left(\sum_{i=1}^{N_{n}}\hat{\mathbb{V}}(\langle S_{n},T_{n}^{(i)}\rangle_{n})-\mathbb{V}(\langle S_{n},T_{n}^{(i)}\rangle_{n}\right)^{2}\right]\longrightarrow 0$.
The proof of this theorem resides in carefully developing for the first part. The convergences are proven by making appear the assumptions 3 and 4 by Cauchy-Schwarz and Jensen inequalities.
Definition 6 (Bona fide estimator).
The bona fide estimators $c$ for the minimization problem 5 are defined, for $i\in\llbracket 1,N_{n}\rrbracket$, as:

### v2

@@ Appendix A Unknown mean formulas
$\hat{\mathbb{V}}\left(\langle S_{n},T_{n}^{(i)}\rangle_{n}\right)=q_{n}^{(0)}\bar{b}_{T^{(i)},n}^{2}+q_{n}^{(1)}\left\langle S_{n}T_{n}^{(i)},T_{n}^{(i)}S_{n}\right\rangle_{n}-q_{n}^{(2)}\langle S_{n},T_{n}^{(i)}\rangle_{n}^{2},$ (10)
with $\text{with }\bar{b}_{T^{(i)},n}^{2}=\frac{p_{n}}{(n-1)^{2}}\sum_{k=1}^{n}\left\langle\tilde{X}_{\cdot k}^{n}\left(\tilde{X}_{\cdot k}^{n}\right)^{T},T_{n}^{(i)}\right\rangle_{n}^{2}-\frac{1}{n}\left\langle S_{n}T_{n}^{(i)},T_{n}^{(i)}S_{n}\right\rangle_{n},$
$\text{and }q_{n}^{(0)}=\frac{(n-1)^{2}}{p_{n}(n-2)(n-3)},q_{n}^{(1)}=\frac{n-1}{p_{n}n(n-2)},q_{n}^{(2)}=\frac{n^{2}-2n-1}{n(n-2)(n-3)}.$
We use these estimators directly in the formulas for $S_{n}^{*}=\mathcal{P}_{\mathcal{S}_{n}^{+}}\left(c_{0}S_{n}+(1-c_{0})\sum_{i=1}^{N_{n}}c_{i}T_{n}^{(i)}\right)$, within the coefficients:
$c_{0}=\min\left((\hat{c}_{0})_{+},1\right),c_{i}=(1-c_{0})\langle S_{n},T_{n}^{(i)}\rangle_{n},$ (11)
with $\text{with }\hat{c}_{0}=$ $\frac{1}{d_{n}^{2}}\Bigg{(}\lVert S_{n}\rVert_{n}^{2}-\hat{\mathbb{V}}(S_{n})-\sum_{k=1}^{N_{n}}\left(\langle S_{n},T_{n}^{(k)}\rangle_{n}^{2}-\hat{\mathbb{V}}\left(\langle S_{n},T_{n}^{(k)}\rangle_{n}\right)\right)\Bigg{)},$
$d_{n}^{2}=$ $\lVert S_{n}\rVert_{n}^{2}-\sum_{k=1}^{N_{n}}\langle S_{n},T_{n}^{(k)}\rangle_{n}^{2}.$
This simple change of definition for $\hat{\mathbb{V}}\left(\langle S_{n},T_{n}^{(i)}\rangle_{n}\right)$ and $\hat{\mathbb{V}}(S_{n})$, depending on the setting of known or unknown mean, conserves the convergence properties we stated in the theoretical part. Every result is proven for both settings.
Appendix B Experimental settings and extensions
In order to improve the complexity of the algorithm, we used the following implementation trick.
@@ Appendix B Experimental settings and extensions
Using this trick, the complexity of our algorithm aligns with the complexity of Single Target Linear Shrinkage, such as the Ledoit-Wolf estimator.
B.1 Experiment 1: Target set impact
$\Sigma$ is built as follows: this is a block diagonal matrix composed of 10 block matrices $(B_{i})_{i=1}^{10}$ of shape $(5,5)$. For each $i\in\llbracket 1,10\rrbracket$, $B_{i}$ is drawn from a Wishart distribution with $5$ degrees of freedom and $\sqrt{11-i}I_{5}$ as scale matrix.
$\{T_{1}^{(i)}\}$ is built as following: the first target $T_{1}^{(1)}=I$, and for $i\in\llbracket 2,10\rrbracket$, the target $T_{1}^{(i)}$ is block-diagonal; where $\Sigma$ has the block $B_{i}$, $T_{1}^{(i)}$ has an Identity block; elsewhere $T_{1}^{(i)}$ has null values.
The set of target is orthogonalized with the Gram-Schmidt algorithm.
For the second set of targets $(T_{2})^{(i)}$, we apply a permutation $\sigma$ on both rows and columns of $T_{1}^{(i)}$, where $\sigma$ is defined as:
B.2 Experiment 3: Heavy-tail robustness
We consider $\Sigma$ block diagonal with $5$ blocks of same size drawn independently from Wishart distributions of $\lfloor p/5\rfloor$ degrees of freedom, and the set of targets is made of $5$ block diagonal matrices, target $i$ with the Identity as block $i$ and 00 elsewhere. We compare, for $20000$ independent experiments using t-Students with degrees of freedom.
Appendix C Theoretical proofs
In this section, we develop the proofs of the theoretical results of the work.
Remark 2.
C.1 Proof of Proposition 1
The minimizer $c^{*}\in\mathbb{R}^{N_{n}+1}$ of:
solves the following equation:
$A_{n}c^{*}=b_{n},$ (17)
with:
$A_{n}=\footnotesize\begin{pmatrix}\lVert S_{n}\rVert_{n}^{2}&\langle S_{n},T_{n}^{(1)}\rangle_{n}&\ldots&\langle S_{n},T_{n}^{(N_{n})}\rangle\\ \langle S_{n},T_{n}^{(1)}\rangle_{n}&1&&(0)\\ \vdots&&\ddots&\\ \langle S_{n},T_{n}^{(N_{n})}\rangle&(0)&&1\end{pmatrix},$ (18)
We have then:
$\det(A_{n})=\lVert S_{n}\rVert_{n}^{2}-\sum_{k=1}^{N_{n}}\langle S_{n},T_{n}^{(k)}\rangle_{n}^{2},$ (19)
$\text{com}(A_{n})^{T}=\footnotesize\begin{pmatrix}1&-\langle S_{n},T_{n}^{(1)}\rangle_{n}&\ldots&-\langle S_{n},T_{n}^{(N)}\rangle_{n}&&\\ -\langle S_{n},T_{n}^{(1)}\rangle_{n}&\\ \vdots&&(B_{n})&\\ -\langle S_{n},T_{n}^{(N)}\rangle_{n}&&&\end{pmatrix},$
with $B_{n}=\left(\lVert S_{n}\rVert_{n}^{2}-\sum_{k=1}^{N_{n}}\langle S_{n},T_{n}^{(k)}\rangle_{n}^{2}\right)I_{N_{n}}+QQ^{T},$ (20)
$Q=\left[\langle S_{n},T_{n}^{(i)}\rangle_{n}\right]_{i=1,...,N_{n}}^{T}.$
Remark 3.
We can conclude by a simple computation:
$c_{0}^{*}=\frac{1}{\det(A_{n})}\left(\langle S_{n},\Sigma_{n}\rangle-\sum_{k=1}^{N_{n}}\langle S_{n},T_{n}^{(k)}\rangle\langle\Sigma_{n},T_{n}^{(k)}\rangle\right),$ (21)
$c_{i}^{*}=\langle\Sigma_{n},T_{n}^{(i)}\rangle-c_{0}^{*}\langle S_{n},T_{n}^{(i)}\rangle.$
C.2 Proof of Lemma 1
We prove firstly an essential lemma on the convergence of the coordinate estimators - of the form $\langle S_{n},T_{n}^{(k)}\rangle_{n}$.
@@ C.2 Proof of Lemma 1
By linearity of the expectation, for all $k\in\llbracket 1,N_{n}\rrbracket$, $\langle S_{n},T_{n}^{(k)}\rangle_{n}$ is un unbiased estimator of $\langle\Sigma_{n},T_{n}^{(k)}\rangle_{n}$.
For the second part of the lemma, let $k\in\llbracket 1,N_{n}\rrbracket$ and let denote $T=T_{n}^{(k)}$. The following proof relies upon the Proposition 1 and its method of graph construction from indices in the Preliminary combinatorial result in [19]. We recall the method and the proposition here.
Let $K\in\mathbb{N}^{*}$, and $K$ indices $(k_{1},...,k_{K})\in\llbracket 1,n\rrbracket^{K}$.
Let’s associate a graph with $K$ vertices $\mathcal{V}=\{1,...,K\}$ to this set of indices. The set of edges $\mathcal{E}$ is built as following: there is an edge between the node $a\in\mathcal{V}$ and $b\in\mathcal{V}$, $a\neq b$ (we don’t allow self-loops), if the corresponding indices are equal, i.e if $k_{a}=k_{b}$. We finally define our graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$.
Proposition 2 (Proposition 1 from [19]).
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ a graph with $K$ vertices generated from some indices $(k_{1}^{(0)},...,k_{K}^{(0)})\in\llbracket 1,n\rrbracket^{K}$ with the procedure described previously. Suppose $\mathcal{G}$ has $C\in\llbracket 1,K\rrbracket$ connected components. Then, there are $\prod_{i=0}^{C-1}(n-i)$ set of indices $(k_{1},...,k_{K})\in\llbracket 1,n\rrbracket^{K}$ which have the associated graph $\mathcal{G}$.
Known mean
When the mean is known, choosing $M^{\prime}\in\mathbb{R}_{+}$ the number of graphs with $4$ nodes and $2$ connected components, we have: